Solve the system by the addition method. Determine whether the equation is independent, dependent, or inconsistent.
-3x+2y=8
3x+2y=8

Answer by xcentaur(357)

-3x+2y=8
3x+2y=8


Adding up we get
(-3x+2y+3x+2y)=8+8
4y=16
y=4


3x=8-2y
3x=8-2(4)
3x=8-8
3x=0
x=0


x=0,y=4


-Prabhat